# On Singular Integrals and Quantitative Rectifiability in Parabolic Space and the Heisenberg Group

### v1

@@ 3.1. The parabolic case
$\alpha_{\mu,k+1}(B(\mathbf{Z}_{E},\ell(E)))<\lambda\text{ or }\alpha_{\mu,k+2}(B(\mathbf{Z}_{E},\ell(E)))\leq\lambda.$
@@ 4. The cylinder blow-up: from qualitative control to quantitative control
The main result of this section is that, if a measure is good for $k$-dimensional Littlewood-Paley kernels, then a Carleson measure condition for dyadic cubes with noticeable $k$-dimensional $\alpha$-numbers implies uniform rectifiability. This is a modification of an analogous argument carried out in the Euclidean setting in [JNT].
@@ 4.1. The set-up
These conditions are very restrictive, but we only have two cases in mind:
In the case of parabolic space, $k\in\{1,...,n-1\}$, and also in this case we will have $\eta\equiv 0$.
In the case of the Heisenberg group, $n=2$, $k=1$, and
@@
Suppose that $\mathbf{X}=(X,t)\in\mathbb{R}^{n+1}$ and $L$ is a $k+2$-dimensional vertical plane. Put $X^{L}=X-X_{L}$, where $X_{L}$ is be the closest point on $\widetilde{L}=\operatorname{Proj}_{\mathbb{R}^{n}}(L)$ to $X\in\mathbb{R}^{n}$ (in the Euclidean metric on $\mathbb{R}^{n}$). Then
@@ 4.2. Main Theorem
$\mathbb{B}_{\lambda,k}(\mu):=\{E\in\mathbb{D}(\mu):\alpha_{k,\mu}(B(\mathbf{Z}_{E},\ell(E))>\lambda\}$
@@
4.3. Upward Domination
@@
We emphasize that we constructed our grid to have the (notationally counterintuitive) property $\operatorname{diam}(E)\leq\ell(E)/4$, so in the integral defining $\beta_{\mu}(E)$ integrates over a ball containing $E$.
@@
Fix $\epsilon>0$. We say that $E^{\prime}\in\mathbb{D}(\mu)$ dominates $E\in\mathbb{D}(\mu)$ from above if $E^{\prime}\supset E$ and
$\beta_{\mu}(E^{\prime})^{2}>\bigg(\frac{\ell(E^{\prime})}{\ell(E)}\bigg)^{\epsilon}\beta_{\mu}(E)^{2}.$
We denote by $\mathbb{D}_{up}(\mu)$ the collection of those dyadic cubes that cannot be dominated from above. So, if $E\in\mathbb{D}_{up}(\mu)$, then for every $E^{\prime}\in\mathbb{D}(\mu)$ with $E^{\prime}\supset E$, we have
(4.3) $\beta_{\mu}(E^{\prime})^{2}\leq\bigg(\frac{\ell(E^{\prime})}{\ell(E)}\bigg)^{\epsilon}\beta_{\mu}(E)^{2}.$
Notice that if $E^{\prime\prime}$ dominates $E^{\prime}$ from above, and $E^{\prime}$ dominates $E$ from above, then $Q^{\prime\prime}$ dominates $Q$ from above.
@@
$\sum_{Q\in\mathbb{D}_{up}(\mu)}\beta_{\mu}^{2}(Q)\mu(E)\geq c(\varepsilon)\sum_{E\in\mathbb{D}(\mu)}\beta_{\mu}^{2}(E)\mu(E).$
Proof.
We first show that if $E\in\mathbb{D}(\mu)\setminus\mathbb{D}_{up}(\mu)$ is such that $\beta(E)\neq 0$, then $E$ can be dominated from above by some $E^{\prime}\in\mathbb{D}_{up}(\mu)$. Indeed, if $E^{\prime}$ dominates $E$ from above, we have
$J(E^{\prime})-J(E)\leq\frac{2}{\epsilon}\log_{2}\bigg(\frac{\beta_{\mu}(E^{\prime})}{\beta_{\mu}(E)}\bigg).$
Consequently, using the fact that $\beta_{\mu}(E_{0})$ is uniformly bounded over $E_{0}\in\mathbb{D}(\mu)$, we see that there are only finitely many $E^{\prime}$ which can dominate $E$ from above. We choose $E^{\prime}$ to be a maximal cube with respect to inclusion which dominates $E$ from above. Then $E^{\prime}\in\mathbb{D}_{up}(\mu)$, for if $E^{\prime}$ is dominated from above by $E^{\prime\prime}$, then $E^{\prime\prime}$ also dominates $E$ from above, which violates the maximality of $E^{\prime}$.
Given $E\in\mathbb{D}(\mu)\setminus\mathbb{D}_{up}(\mu)$, let $\widetilde{E}$ by the maximal cube of $\mathbb{D}_{up}(\mu)$ which dominates $E$ from above. Fix some $E_{0}\in\mathbb{D}_{up}(\mu)$. Then
$\sum_{\begin{subarray}{c}E\in\mathbb{D}(\mu)\setminus\mathbb{D}_{up}(\mu)\\ \widetilde{E}=E_{0}\end{subarray}}\beta_{\mu}(E)^{2}\mu(E)=$ $\sum_{m\geq 1}\sum_{\begin{subarray}{c}E\in\mathbb{D}(\mu)\setminus\mathbb{D}_{up}(\mu)\\ \ell(E)=2^{-m}\ell(E_{0}),\,\widetilde{E}=E_{0}\end{subarray}}\beta_{\mu}(E)^{2}\mu(E)$
$\leq$ $\sum_{m\geq 1}2^{-\epsilon m}\beta_{\mu}(E_{0})^{2}\bigg[\sum_{\begin{subarray}{c}E\in\mathbb{D},E\subset E_{0}\\ \ell(E)=2^{-m}\ell(E_{0})\end{subarray}}\mu(E)\bigg]$
$\leq$ $C(\epsilon)\beta_{\mu}(E_{0})^{2}\mu(E_{0})$
The conclusion of the lemma follows from summing over all $E_{0}\in\mathbb{D}_{up}(\mu)$. ∎
@@ 4.4. Planar concentration via square function coefficients
We fix a function $\varphi(X,t)=X\psi(X,t)$ where $\psi:\mathbb{R}^{n+1}\rightarrow\mathbb{R}$ is Lipschitz with $\|\psi\|_{lip}\lesssim 1$, supported in $B_{9}(\mathbf{0})$, and equal to $1$ on $B_{6}(\mathbf{0})$.
@@
$\int_{B_{2r}(\mathbf{Z})\setminus H_{3\beta}}\left(\frac{\operatorname{dist}(\mathbf{X},H)}{r}\right)^{2}d\mu(\mathbf{X})\leq C\Theta_{\varphi,\mu,2}(\mathbf{X},r).$
@@
Let $\Delta>0$, $r>0$, and $\mathbf{Z}\in\operatorname{supp}(\mu)$. Suppose that, for some vertical $k$-plane $L$, we have
@@ 4.5. The Cylindrical Blow-Up
There exists $\Lambda>1$, $\Delta>0$, $\rho>0$, and $N\in\mathbb{N}$, along with a family $\mathcal{F}\subset\operatorname{Lip}_{0}^{\operatorname{odd}}(B(\mathbf{0},\Lambda))$ such that for every dyadic cube $E\in\mathbb{D}_{up}(\mu)$, if
$\alpha_{k,\mu}(\mathbf{Z}_{E},\Lambda\cdot\ell(E))\leq\rho,$
@@
(4.5) $\max_{\varphi\in\mathcal{F}}\Theta_{\mu,\varphi,2\Lambda}(E)\geq\Delta\beta_{\mu}(E)^{2}\ell(E)^{k+2},$
where $\mathcal{F}\leq C(N)$ and $\max_{\varphi\in\mathcal{F}}\|\varphi\|_{lip}\leq C(N)$.
@@
Fix $\Delta$, $\Lambda$, $\rho$ and $\mathcal{F}$ as in the proof of Proposition 4.8. Put $\mathcal{Q}_{\rho}=\{E\subset\mathbb{D}(\mu):\,\alpha_{k,\mu}(B(\mathbf{Z}_{E},\Lambda\cdot\ell(E))>\rho\}$. Reasoning in an analogous way to how Lemma 2.5 follows from Lemma 2.4, we see that for each cube $E_{0}\in\mathbb{D}(\mu)$,
(4.6) $\Delta\sum_{\begin{subarray}{c}E\in\mathbb{D}_{\operatorname{up}}(\mu)\backslash\mathcal{Q}_{\rho}:\\ E\subset E_{0}\end{subarray}}\beta_{\mu}(\mathbf{Z}_{E},\ell(E))^{2}\mu(E)\lesssim_{\mu,\mathcal{F}}\mu(E_{0}).$
But since, by assumption and Lemma 2.6, $\mathcal{Q}_{\rho}$ is Carleson, and $\beta_{\mu}(B(\mathbf{Z}_{E},\ell(E)))\leq 1$,
$\sum_{\begin{subarray}{c}E\in\mathbb{D}_{\operatorname{up}}(\mu):\\ E\subset E_{0}\end{subarray}}\beta_{\mu}(\mathbf{Z}_{E},\ell(E))^{2}\mu(E)\lesssim\ell(E_{0})^{k+2},$
where the implicit constant now can depend on the Carleson norm of $\mathcal{Q}_{\rho}$ in addition to $\Delta$, $A$, and $\mathcal{F}$. Finally, Lemma 4.5 completes the proof of the discrete uniform rectifiability property.∎
In preparation for the proof of Proposition 4.8, we will require some Lipschitz functions:
Choose $\psi_{0}\in\operatorname{Lip}_{0}(B(\mathbf{0},\Lambda))$ with $\psi_{0}\equiv 1\text{ on }B(\mathbf{0},3\Lambda/4)$ and $\psi_{0}\lesssim 1$. Select a countable dense set $\{\psi_{j}\}_{j\geq 0}$ in the collection
$\operatorname{Lip}_{0}^{\operatorname{even}}(B(\mathbf{0},\Lambda))=\big\{\psi\in\operatorname{Lip}_{0}(B(\mathbf{0},\Lambda)):\psi(X,t)=\psi(-X,t)\text{ for all }(X,t)\in B(\mathbf{0},\Lambda)\bigl\}$
$\text{ satisfying }\|\psi\|_{\operatorname{Lip}}\leq 1.$
Then we consider the functions $\widetilde{\psi}(\mathbf{X})=X\psi(\mathbf{X},t)$. For each $j\geq 0$, $\widetilde{\psi}_{j}\in\operatorname{Lip}_{0}^{\operatorname{odd}}(B(\mathbf{0},\Lambda))$, with [^7] The product of two bounded Lipschitz functions is a Lipschitz function. $\|\widetilde{\psi}_{0}\|_{\operatorname{Lip}}\lesssim_{\Lambda}1$. We will set $\mathcal{F}_{\ell}=\{\widetilde{\psi}_{0},\dots,\widetilde{\psi}_{\ell}\}$.
@@ 4.5.1. The Limit Measure
$\max_{\varphi\in\mathcal{F}_{l}}\Theta_{\widetilde{\mu}_{l},\varphi,\ell}(\widetilde{E}_{\ell})\leq\frac{1}{l}\beta^{2}_{\widetilde{\mu}_{l}}(\widetilde{E}_{l})\ell(E_{\ell})^{k+2}.$
We rescale and re-normalize the measure $\widetilde{\mu}_{l}$ to a $(k+2)$-ADR measure $\mu_{l}$ and such that $\widetilde{E}_{l}$ gets mapped to cube $E_{l}$ with $\ell(E_{l})=1$ and $\mathbf{Z}_{\widetilde{E}_{l}}$ is mapped to the origin $\mathbf{0}$. This rescaling and renormalization is applied to the grid $\mathbb{D}(\widetilde{\mu}_{l})$ to produce a new grid $\mathbb{D}(\mu_{l})$. The image of $\widetilde{E}_{l}\in\mathbb{D}(\widetilde{\mu}_{l})$ is $E_{l}\in\mathbb{D}(\mu_{l})$ where $\mathbf{Z}_{E_{\ell}}=\mathbf{0}$. We have
(4.7) $\max_{\varphi\in\mathcal{F}_{l}}\Theta_{\mu_{l},\varphi,l}(E_{l})\leq\frac{1}{l}\beta^{2}_{\mu_{l}}(E_{l})=\frac{1}{l}\beta_{\mu_{l}}^{2}(\mathbf{0},1).$
Because the measures $\mu_{l}$ enjoy the ADR property with uniform bounds on the ADR constants, we can pass to a subsequence which converges to some measure $\mu$ such that
@@
Since $\alpha_{k,\mu_{\ell}}(\mathbf{0},\Lambda)\leq\frac{1}{\ell}$, we have that $\alpha_{\mu}(\mathbf{0},\Lambda)=0$ (Lemma 2.7). Therefore, $\operatorname{supp}(\mu)\cap B(\mathbf{0},\Lambda)$ is contained in a vertical plane, and therefore $\beta_{\mu}(\mathbf{0},1)=0.$ Therefore we may assume that $\lim_{l\rightarrow\infty}\beta_{\mu_{l}}(\mathbf{0},1)=0$.
Since $E_{l}\in\mathbb{D}_{up}(\mu_{l})$, we next observe that
(4.8) $\beta_{\mu_{l}}(\mathbf{0},\Lambda)\lesssim\Lambda^{\varepsilon}\beta_{\mu_{l}}(\mathbf{0},1).$
This is a consequence of the definition of $\mathbb{D}_{up}(\mu)$, and one can verify it by applying the condition to a dyadic cube $E\in\mathbb{D}_{\operatorname{up}}(\mu)$ containing the origin such that $\ell(E)\approx\Lambda$.
Our ultimate goal will be to show that (4.8) fails to hold for large $\ell$ if $\Lambda$ is chosen appropriately. The selection of $\Lambda$ is not particularly subtle, and, say $\Lambda=3000$ would certainly do the job, but we carry it around as a parameter.
Set $\beta_{l}=\beta_{\mu_{l}}(\mathbf{0},\Lambda)$. Let $L_{l}$ be an optimizing plane for $\beta_{\mu_{l}}(\mathbf{0},\Lambda)$. By passing to a subsequence, if necessary, we may assume that the planes $L_{l}$ converge to a plane $L$ locally in the Hausdorff metric. As a consequence of Corollary 4.7, there is a constant $C_{1}>0$ such that
(4.9) $\int_{B_{\Lambda/5}(\mathbf{0})\setminus L_{l,C_{1}\beta_{l}}}\bigg(\frac{\operatorname{dist}(\mathbf{Y},L_{l})}{\beta_{l}}\bigg)^{2}d\mu_{l}(\mathbf{Y})\lesssim\frac{1}{\ell}.$
The stretched measure $\nu$:. Let $\mathcal{R}^{l}$ be the composition of translation and horizontal rotation (with $\mathcal{O}^{l}$ denoting the horizontal rotation) such that $\mathcal{R}^{l}(L_{l})=\{0_{n-k}\}\times\mathbb{R}^{k}\times\mathbb{R}$, where $0_{n-k}$ is the origin of $\in\mathbb{R}^{n-k}$. We can pass to another subsequence to ensure that $\mathcal{O}^{l}$ converges to a horizontal rotation $\mathcal{O}$. Since $\beta_{\ell}\to 0$ as $\ell\to\infty$ and $0\in\operatorname{supp}(\mu_{\ell})$, we also have $\mathcal{R}^{\ell}\to\mathcal{O}$ as $\ell\to\infty$. We introduce (spatial) coordinates $X=(X^{\prime},X^{\prime\prime})\in\mathbb{R}^{n-k}\times\mathbb{R}^{k}$.
If $\ell$ is sufficiently large, the plane $L_{l}$ goes very close to to the origin (recall (4.9)), and since horizontal rotations are isometries we see that for all large $\ell$, $(\mathcal{R}^{\ell})^{-1}B(0,\Lambda/6)\subset B(0,\Lambda/5)$. Therefore, recalling Lemma 4.2, we deduce for all sufficiently large $\ell$ that
(4.10) $\int_{B(\mathbf{0},\Lambda/6)\backslash([-C_{1}\beta_{\ell},C_{1}\beta_{\ell}]^{n-k}\times\mathbb{R}^{k}\times\mathbb{R})}\Bigl(\frac{|Y^{\prime}|}{\beta_{\ell}}\Bigl)^{2}d\mu_{\ell}((\mathcal{R}^{\ell})^{-1}(\mathbf{Y}))\lesssim\frac{1}{\ell}.$
@@
(4.11) $\nu_{l}=\mu_{l}((\mathcal{R}^{l})^{-1}\circ\mathcal{S}_{C_{1}\beta_{l}}(\,\cdot\,)).$

### v2

@@ 3.1. The parabolic case
$\alpha_{\mu,k+1}(\mathbf{Z}_{E},\ell(E))<\lambda\text{ or }\alpha_{\mu,k+2}(\mathbf{Z}_{E},\ell(E))\leq\lambda.$
@@ 4. The cylinder blow-up: from qualitative control to quantitative control
The main result of this section is that, if a measure is good for $k$-dimensional Littlewood-Paley kernels, then a Carleson measure condition for dyadic cubes with noticeable $k$-dimensional $\alpha$-numbers implies uniform rectifiability. This is a modification of an analogous argument carried out in the Euclidean setting in [35].
@@ 4.1. The set-up
These conditions are very restrictive, but we only have two cases in mind. In the case of parabolic space, $k\in\{1,...,n-1\}$, and also in this case we will have $\eta\equiv 0$. In the case of the Heisenberg group, $n=2$, $k=1$, and
@@
Suppose that $\mathbf{X}=(X,t)\in\mathbb{R}^{n+1}$ and $L$ is a $(k+2)$-dimensional vertical plane. Put $X^{L}=X-X_{L}$, where $X_{L}$ is be the closest point on $\widetilde{L}=\operatorname{Proj}_{\mathbb{R}^{n}}(L)$ to $X\in\mathbb{R}^{n}$ (in the Euclidean metric on $\mathbb{R}^{n}$). Then
@@ 4.2. Main Theorem
$\mathbb{B}_{\lambda,k}(\mu):=\{E\in\mathbb{D}(\mu):\alpha_{k,\mu}(\mathbf{Z}_{E},\ell(E))>\lambda\}$
@@
For the remainder of section 4 we will fix a measure $\mu$ satisfying the hypotheses of Theorem 4.3.
4.3. Doubling $\beta$-number
We fix a parameter $\Lambda$, which we assume is a power of two. Its choices will not be too subtle and can be fixed at this point, but we leave it free for now.
@@ 4.3. Doubling $\beta$-number
We emphasize that we constructed our grid to have the (notationally counterintuitive) property $\operatorname{diam}(E)\leq\ell(E)/4$, so in the integral defining $\beta_{\nu}(E)$ integrates over a ball containing $E$.
@@
We say that $E\in\mathbb{D}(\mu)$ is $\beta$-doubling if
$\beta_{\mu}(\mathbf{Z}_{E},\Lambda\ell(E)))\leq\Lambda\beta_{\mu}(E).$
We denote by $\mathbb{D}_{up}(\mu)$ the collection of $\beta$-doubling dyadic cubes.
Our goal is to prove the following lemma.
@@
There is a constant $C(\Lambda)$ such that for every $E_{0}\in\mathbb{D}(\mu)$
$\sum_{\begin{subarray}{c}E\in\mathbb{D}(\mu)\\ E\subset E_{0}\end{subarray}}\beta_{\mu}^{2}(E)\mu(E)\leq C(\Lambda)\Bigl(\sum_{\begin{subarray}{c}E\in\mathbb{D}_{up}(\mu)\\ E\subset E_{0}\end{subarray}}\beta_{\mu}^{2}(E)\mu(E)+\mu(E_{0})\Bigl).$
In verifying this lemma, it will be convenient, for $E\in\mathbb{D}(\mu)$, to denote by $E_{\Lambda}$ the dyadic cube $E_{\Lambda}\in\mathbb{D}(\mu)$ with $\ell(E_{\Lambda})=2\Lambda\ell(E)$. Since $B_{\ell(E_{\Lambda})}(\mathbf{Z}_{E_{\Lambda}})\supset B_{\Lambda\ell(E)}(\mathbf{Z}_{E})$, if $E$ is not $\beta$-doubling, then there is an absolute constant $C>0$ such that
(4.3) $\beta_{\mu}(E)\leq\frac{1}{\Lambda}\beta_{\mu}(\mathbf{Z}_{E},3\Lambda\ell(E))\leq\frac{C\ell(E)}{\ell(E_{\Lambda})}\beta_{\mu}(E_{\Lambda})\leq\Bigl(\frac{\ell(E)}{\ell(E_{\Lambda})}\Bigl)^{c}\beta_{\mu},(E_{\Lambda}).$
and therefore, provided $\Lambda$ is chosen appropriately, there is a constant $c>0$ such that
(4.4) $\beta_{\mu}(E)\leq\Bigl(\frac{\ell(E)}{\ell(E_{\Lambda})}\Bigl)^{c}\beta_{\mu}(E_{\Lambda}).$
We call $E^{\prime}$ a $\Lambda$-ancestor of $E$ if $E^{\prime}\supset E$, $\ell(E^{\prime})/\ell(E)=(2\Lambda)^{d}$ for a non-negative integer $d$.
Suppose $E\in\mathbb{D}(\mu)\backslash\mathbb{D}_{\operatorname{up}}(\mu)$. If $\beta_{\mu}(E)>0$, then since also $\beta_{\mu}(E^{\prime})\lesssim 1$ for every $E^{\prime}\in\mathbb{D}(\mu),$ there is a smallest $\Lambda$-ancestor $E_{d}$ of $E$ with $E_{d}\in\mathbb{D}_{\operatorname{up}}(\mu)$. By iterating (4.4) we have
$\beta_{\mu}(E)\leq\Bigl(\frac{\ell(E)}{\ell(E^{\prime})}\Bigl)^{c}\beta_{\mu}(E^{\prime})$
whenever $E^{\prime}$ is a $\Lambda$-dyadic ancestor of $E$ and $E^{\prime}\subset E_{d}$. But, now if $E^{\prime}$ is any dyadic ancestor of $E$ with $E\subset E^{\prime}\subset E_{d}$, then by choose $E^{\prime\prime}$ to be the largest $\Lambda$-dyadic ancestor of $E$ with $E^{\prime\prime}\subset E^{\prime}$, we see that $\beta_{\mu}(E^{\prime\prime})\lesssim_{\Lambda}\beta_{\mu}(E^{\prime})$, and therefore
(4.5) $\beta_{\mu}(E)\lesssim_{\Lambda}\Bigl(\frac{\ell(E)}{\ell(E^{\prime})}\Bigl)^{c}\beta_{\mu}(E^{\prime})\text{ whenever }E^{\prime}\in\mathbb{D}(\mu)\text{ satisfies }E\subset E^{\prime}\subset E_{d}.$
[…]
$\sum_{\begin{subarray}{c}E\in\mathbb{D}(\mu)\setminus\mathbb{D}_{up}(\mu)\\ \widetilde{E}=F\end{subarray}}\beta_{\mu}(E)^{2}\mu(E)=$ $\sum_{m\geq 1}\sum_{\begin{subarray}{c}E\in\mathbb{D}(\mu)\setminus\mathbb{D}_{up}(\mu)\\ \ell(E)=2^{-m}\ell(F),\,\widetilde{E}=F\end{subarray}}\beta_{\mu}(E)^{2}\mu(E)$
$\lesssim_{\Lambda}$ $\sum_{m\geq 1}2^{-cm}\beta_{\mu}(F)^{2}\bigg[\sum_{\begin{subarray}{c}E\in\mathbb{D}(\mu),E\subset F\\ \ell(E)=2^{-m}\ell(F)\end{subarray}}\mu(E)\bigg]$
$\leq_{\Lambda}$ $\beta_{\mu}(F)^{2}\mu(F)$
The conclusion of the lemma follows from summing over all possible $F$.
@@ 4.4. Planar concentration via square function coefficients
We fix a function $\varphi(X,t)=X\psi(X,t)$ where $\psi:\mathbb{R}^{n+1}\rightarrow\mathbb{R}$ is Lipschitz with $\|\psi\|_{\operatorname{Lip}}\lesssim 1$, supported in $B_{9}(\mathbf{0})$, and equal to $1$ on $B_{6}(\mathbf{0})$.
@@
$\int_{B_{2r}(\mathbf{Z})\setminus H_{3\beta}}\left(\frac{\operatorname{dist}(\mathbf{X},H)}{r}\right)^{2}d\mu(\mathbf{X})\leq C\Theta_{\varphi,\mu,2}(\mathbf{Z},r).$
@@
Let $\Delta>0$, $r>0$, and $\mathbf{Z}\in\operatorname{supp}(\nu)$. Suppose that, for some vertical $k$-plane $L$, we have
@@ 4.5. The Cylindrical Blow-Up
There exists $\Delta>0$, $\rho>0$, and $N\in\mathbb{N}$, along with a family $\mathcal{F}\subset\operatorname{Lip}_{0}^{\operatorname{odd}}(B_{\Lambda}(\mathbf{0}))$, where $\mathcal{F}\leq C(N)$ and $\max_{\varphi\in\mathcal{F}}\|\varphi\|_{\operatorname{Lip}}\leq C(N)$, such that the following holds. Let $\Lambda\geq 2^{12}=4096$ be a power of $2$. Then, for every dyadic cube $E\in\mathbb{D}_{up}(\mu)$, if
$\alpha_{k,\mu}(\mathbf{Z}_{E},\Lambda\ell(E))\leq\rho,$
@@
(4.7) $\max_{\varphi\in\mathcal{F}}\Theta_{\mu,\varphi,2\Lambda}(E)\geq\Delta\beta_{\mu}(E)^{2}\ell(E)^{k+2}.$
@@
Pick a cube $E_{0}\in\mathbb{D}(\mu)$ and put
$\mathcal{Q}_{\rho}=\{E\subset\mathbb{D}(\mu):\,\alpha_{k,\mu}(\mathbf{Z}_{E},\Lambda\ell(E))>\rho\}.$
Reasoning in an analogous way to how Lemma 2.5 follows from Lemma 2.4, we see that
(4.8) $\Delta\sum_{E\in\mathbb{D}_{\operatorname{up}}(\mu)\backslash\mathcal{Q}_{\rho}}\beta_{\mu}(\mathbf{Z}_{E},\ell(E))^{2}\mu(E)\lesssim_{\mu,\mathcal{F}}\mu(E_{0}).$
But since $\mathcal{Q}_{\rho}$ is Carleson (by hypothesis and Lemma 2.6), and $\beta_{\mu}(\mathbf{Z}_{E},\ell(E))\lesssim 1$,
$\sum_{E\in\mathbb{D}_{\operatorname{up}}(\mu)}\beta_{\mu}(\mathbf{Z}_{E},\ell(E))^{2}\mu(E)\lesssim\ell(E_{0})^{k+2},$
where the implicit constant now can depend on the Carleson norm of $\mathcal{Q}_{\rho}$ in addition to $\Delta$, $A$, and $\mathcal{F}$. Finally, an application of Lemma 4.5 gives
$\sum_{\begin{subarray}{c}E\in\mathbb{D}(\mu)\\ E\subset E_{0}\end{subarray}}\beta_{\mu}(\mathbf{Z}_{E},\ell(E))^{2}\mu(E)\lesssim\ell(E_{0})^{k+2}.$
In preparation for the proof of Proposition 4.8, we will require some Lipschitz functions. We will fix $\Lambda>1$, which will be taken to be large. Choose $\psi_{0}\in\operatorname{Lip}_{0}(B(\mathbf{0},\Lambda))$ with $\psi_{0}\equiv 1\text{ on }B(\mathbf{0},3\Lambda/4)$ and $\psi_{0}\lesssim 1$. Select a countable dense set $\{\psi_{j}\}_{j\geq 0}$ in the collection (we are forcing $\psi_{0}$ to be the function described above)
$\operatorname{Lip}_{0}^{\operatorname{even}}(B(\mathbf{0},\Lambda))\\ =\big\{\psi\in\operatorname{Lip}_{0}(B(\mathbf{0},\Lambda)):\psi(X,t)=\psi(-X,t)\text{ for all }(X,t)\in B(\mathbf{0},\Lambda)\bigl\}\cap\{\|\psi\|_{\operatorname{Lip}}\leq 1\}.$
Then we consider the functions $\widetilde{\psi}(\mathbf{X})=X\psi(\mathbf{X})$. For each $j\geq 0$, $\widetilde{\psi}_{j}\in\operatorname{Lip}_{0}^{\operatorname{odd}}(B(\mathbf{0},\Lambda))$, with [^8] The product of two bounded Lipschitz functions is a Lipschitz function. $\|\widetilde{\psi}_{0}\|_{\operatorname{Lip}}\lesssim_{\Lambda}1$. We will set $\mathcal{F}_{\ell}=\{\widetilde{\psi}_{0},\dots,\widetilde{\psi}_{\ell}\}$.
@@ 4.5.1. The Limit Measure
$\max_{\varphi\in\mathcal{F}_{\ell}}\Theta_{\widetilde{\mu}_{\ell},\varphi,\ell}(\widetilde{E}_{\ell})\leq\frac{1}{\ell}\beta^{2}_{\widetilde{\mu}_{\ell}}(\widetilde{E}_{\ell})\ell(E_{\ell})^{k+2}.$
@@
(4.9) $\max_{\varphi\in\mathcal{F}_{\ell}}\Theta_{\mu_{\ell},\varphi,\ell}(E_{\ell})\leq\frac{1}{\ell}\beta^{2}_{\mu_{\ell}}(E_{\ell})=\frac{1}{\ell}\beta_{\mu_{\ell}}^{2}(\mathbf{0},1).$
Because the measures $\mu_{\ell}$ enjoy the ADR property with uniform bounds on the ADR constants, we can pass to a subsequence which converges to some measure $\mu$ such that
[…]
Since $\alpha_{k,\mu_{\ell}}(\mathbf{0},\Lambda)\leq\frac{1}{\ell}$, we have that $\alpha_{\mu}(\mathbf{0},\Lambda)=0$ (Lemma 2.7). Therefore, $\operatorname{supp}(\mu)\cap B(\mathbf{0},\Lambda)$ is contained in a vertical $k$-plane, and therefore $\beta_{\mu}(\mathbf{0},1)=0.$ Thus, we may assume that $\lim_{l\rightarrow\infty}\beta_{\mu_{\ell}}(\mathbf{0},1)=0$.
Since $E_{\ell}\in\mathbb{D}_{up}(\mu_{\ell})$, we next observe that, for some $\epsilon>0$
(4.10) $\beta_{\mu_{\ell}}(\mathbf{0},\Lambda)\lesssim\Lambda\beta_{\mu_{\ell}}(\mathbf{0},1).$
Our ultimate goal will be to show that (4.10) fails to hold for large $\ell$.
Set $\beta_{\ell}=\beta_{\mu_{\ell}}(\mathbf{0},\Lambda)$. Let $L_{\ell}$ be an optimizing plane for $\beta_{\mu_{\ell}}(\mathbf{0},\Lambda)$. By passing to a subsequence, if necessary, we may assume that the planes $L_{\ell}$ converge to a plane $L$ locally in the Hausdorff metric. As a consequence of Corollary 4.7, there is a constant $C_{1}>0$ such that
(4.11) $\int_{B_{\Lambda/5}(\mathbf{0})\setminus L_{\ell,C_{1}\beta_{\ell}}}\bigg(\frac{\operatorname{dist}(\mathbf{Y},L_{\ell})}{\beta_{\ell}}\bigg)^{2}d\mu_{\ell}(\mathbf{Y})\lesssim\frac{1}{\ell}.$
The stretched measure $\sigma$: Let $\mathcal{R}^{\ell}$ be the composition of translation and horizontal rotation (with $\mathcal{O}^{\ell}$ denoting the horizontal rotation) such that $\mathcal{R}^{\ell}(L_{\ell})=\{0_{n-k}\}\times\mathbb{R}^{k}\times\mathbb{R}$, where $0_{n-k}$ is the origin of $\in\mathbb{R}^{n-k}$. We can pass to another subsequence to ensure that $\mathcal{O}^{\ell}$ converges to a horizontal rotation $\mathcal{O}$. Since $\beta_{\ell}\to 0$ as $\ell\to\infty$ and $0\in\operatorname{supp}(\mu_{\ell})$, we also have $\mathcal{R}^{\ell}\to\mathcal{O}$ as $\ell\to\infty$. We introduce (spatial) coordinates $X=(X^{\prime},X^{\prime\prime})\in\mathbb{R}^{n-k}\times\mathbb{R}^{k}$.
If $\ell$ is sufficiently large, the plane $L_{\ell}$ goes very close to to the origin (recall (4.11)), and since horizontal rotations are isometries we see that for all large $\ell$, $(\mathcal{R}^{\ell})^{-1}B_{\Lambda/6}(\mathbf{0})\subset B_{\Lambda/5}(\mathbf{0})$. Therefore, recalling Lemma 4.2, we deduce for all sufficiently large $\ell$ that
(4.12) $\int_{B_{\Lambda/6}(\mathbf{0})\backslash([-C_{1}\beta_{\ell},C_{1}\beta_{\ell}]^{n-k}\times\mathbb{R}^{k}\times\mathbb{R})}\Bigl(\frac{|Y^{\prime}|}{\beta_{\ell}}\Bigl)^{2}d\mu_{\ell}((\mathcal{R}^{\ell})^{-1}(\mathbf{Y}))\lesssim\frac{1}{\ell}.$
@@
(4.13) $\sigma_{\ell}=\mu_{\ell}((\mathcal{R}^{\ell})^{-1}\circ\mathcal{S}_{C_{1}\beta_{\ell}}(\,\cdot\,)).$